# Resonance in black hole ringdown: Benchmarking quasinormal mode excitation and extraction

### v1

@@ Abstract
We investigate how resonant excitation near exceptional points manifests in Kerr black hole ringdown waveforms and examine its extraction. Using waveforms generated by localized initial data, where quasinormal mode amplitudes are given solely by excitation factors, we establish a controlled benchmark for overtone extraction. Applying an iterative fitting method with mirror modes, we analyze a mild resonance in the $(l,m)=(2,2)$ multipole and a sharp resonance in the $(3,1)$ multipole occurring as part of a sequence of successive resonances. For $(2,2)$, we extract the fundamental mode, the first three overtones, and the fundamental mirror mode with relative errors below $10\%$, and show that residual waveforms exhibit the expected damped sinusoids together with distinctive resonance signatures. For $(3,1)$, we demonstrate that resonances can not only amplify but also reduce QNM excitations, reshaping the overtone hierarchy and rendering the sharp resonance more pronounced in ringdown. Our results clarify the imprint of resonance in Kerr ringdown and highlight both the robustness and limitations of current extraction techniques, providing a foundation for more reliable extraction of higher overtones and for applications to observational data analysis.
@@ I Introduction
The ringdown phase of gravitational waves emitted by perturbed black holes encodes fundamental information about the spacetime geometry in the strong-field regime (see reviews [1, 2, 3]). In this phase, the signal is well described by a superposition of quasinormal modes (QNMs), whose complex frequencies are determined solely by the background spacetime. In addition to the QNM frequencies, their complex amplitudes also carry key information about the strong-field physics, such as the excitation process and the spacetime structure. The QNM amplitudes can be factorized into a term dependent on the initial data and an excitation factor, the latter of which depends only on the background spacetime.
@@
A particularly intriguing feature arises when QNM frequencies approach each other in the complex plane. This leads to an avoided crossing in which the QNM frequencies trace a hyperbolic trajectory, together with resonant amplification of the corresponding excitation factors in a lemniscate pattern [10, 11], which has been proposed recently as a resolution of the longstanding mystery of the “dissonance” anomaly in Kerr QNM spectrum [12], as well as a novel diagnostic of black hole structure and a potential probe of new physics. This phenomenon universally occurs near exceptional points [13]—a characteristic feature of non-Hermitian systems. Similar behavior has also been reported in Einstein-Maxwell-axion system [14], reinforcing the potential of the resonance to probe new physics.
@@
The aim of the present work is to identify and characterize the features of resonant excitation in the Kerr ringdown waveform. Our analysis combines three key ingredients: recently computed excitation factors [10, 11], the iterative extraction method [26], and a Kerr ringdown waveform generated by localized initial data [4]. In this waveform, modeled as a delta-function source term, the QNM amplitudes are given solely by the excitation factors. This setup provides a controlled and idealized environment for exploring resonance features and testing their extraction from a waveform composed of an infinite QNM sum and power-law tails. While the response of a localized source has been studied in previous work [4, 30, 15], to our knowledge this is the first time that the fitting of the resulting waveform is explicitly performed. Since QNM amplitudes in more realistic waveforms are also largely governed by excitation factors, the insights gained from this analysis can be directly applied to the fitting of observational signals.
@@ II.1 Teukolsky formalism
Gravitational waves emitted in the ringdown phase of binary black hole mergers are described by linear perturbations on a background black hole spacetime (see [31, 32] for reviews). The perturbations on the Kerr spacetime are governed by a master equation called the Teukolsky equation [33, 34, 35] derived within the Newman-Penrose formalism [36]. The master variable for the Teukolsky equation is $\rho^{-4}\Psi_{4}$, where $\rho=(r-ia\cos\theta)^{-1}$ and $\Psi_{4}$ is the component of the Weyl tensor projected onto null tetrads in the Newman-Penrose formalism. In the asymptotic region far from perturbed Kerr black hole, $\Psi_{4}$ is related to the gravitational wave strain by $\Psi_{4}=\frac{1}{8}(\partial_{t}-\partial_{r})^{2}(h_{+}-ih_{\times})$. Decomposing the master variable
@@
$\int_{0}^{\pi}S^{*}_{lm}(\theta;c)S_{lm}(\theta;c)\sin\theta\mathrm{d}\theta=1.$ (4)
The radial Green’s function is constructed from two linearly independent homogeneous solutions: the ‘in’ solution $R^{\text{in}}_{lm}$ satisfying the boundary condition
@@
and the ‘up’ solution $R^{\text{up}}_{lm}$ satisfying
@@
where $B^{\text{inc}}_{lm}$, $B^{\text{ref}}_{lm}$, $C^{\text{inc}}_{lm}$, and $C^{\text{ref}}_{lm}$ are complex asymptotic amplitudes and $k=\omega-am/(2Mr_{+})$. $r_{*}$ is the tortoise coordinate which satisfies $\mathrm{d}r_{*}/\mathrm{d}r=(r^{2}+a^{2})/\Delta$. We fix the integration constant and employ the conventional tortoise coordinate:
@@
$\Psi_{4}(t,r\to\infty,\theta,\phi)=\frac{1}{2\pi}\frac{1}{r}\sum_{l=2}^{\infty}\sum_{m=-l}^{l}\int\mathrm{d}\omega\frac{B^{\text{ref}}_{lm}}{2i\omega B^{\text{inc}}_{lm}}I_{lm}S_{lm}(\theta;a\omega)\frac{e^{im\varphi}}{\sqrt{2\pi}}e^{-i\omega(t-r_{*})}$ (13)
@@ II.2 QNM decomposition
While Eq. (13) involves an integration along the real frequency axis, by enclosing the integration contour on the complex frequency plane, we can rewrite it as a superposition of QNMs together with a power-law tail, provided that no pole arises in the source-dependent term. By evaluating the complex frequency integral using the residue theorem, we obtain
@@
where $B_{lmn}$ is the excitation factor defined by
@@
(a) QNM frequency
(b) Excitation factor
Figure 1: QNM frequencies (left) and excitation factors (right) for the $(l,m)=(2,2)$ mode with spin parameter $0\leq a/M\leq 1-10^{-6}$. The triangle marker in the left panel denotes the Schwarzschild value. The range $0.85\leq a/M\leq 0.95$ is highlighted with thick curves, and the arrow in the right panel shows the direction of increasing spin. All figures involving QNM frequencies or excitation factors use the Reference Dataset (RD) [10, 11].
@@
Throughout this paper, we adopt a convention where positive overtone indices $n=+0,+1,+2,\cdots$ correspond to original modes, and negative indices $n=-0,-1,-2,\cdots$ to mirror modes. Note that we explicitly include $\pm 0$ to distinguish the original fundamental mode from its mirror counterpart.
@@
There exists a useful symmetry relation between the QNM frequencies and excitation factors of the original and mirror modes. The homogeneous radial Teukolsky equation is invariant under the combined operation of complex conjugation and the replacements $m\to-m$ and $\omega\to-\omega^{*}$. This symmetry implies that the mirror mode frequency is related to the original mode frequency via
@@
Since the same symmetry applies to the radial Teukolsky solutions, the ‘in’ solution satisfies $R^{\text{in*}}_{lm}(\omega)=R^{\text{in}}_{l-m}(-\omega^{*})$. Consequently, the asymptotic amplitudes also satisfy a similar relation, $B^{\text{ref/inc*}}_{lm}(\omega)=B^{\text{ref/inc}}_{l-m}(-\omega^{*})$. Using this relation with Eqs. (16) and (18), the excitation factor of the mirror mode is related to that of the original mode via 222 Equation (19) exactly matches Eq. (39) of Ref. [45] aside from notation, while it is inconsistent with Eq. (7) of Ref. [15], which can only be obtained by assuming the opposite sign on the right-hand side of Eq. (19).
@@
Figure 1 shows the QNM frequencies and excitation factors for $(l,m)=(2,2)$ up to the seventh overtone. We highlight the region with the spin parameter $0.85\leq a/M\leq 0.95$ using thick curves. As seen in Fig. 1, the excitation factors of the fifth and sixth overtones exhibit a significant enhancement around $a/M=0.9$, where the corresponding QNM frequencies approach each other. This resonant excitation phenomenon universally occurs in the QNMs in the vicinity of exceptional points [10]. Near this spin value, the excitation factors exhibit a distinctive knot-shaped loop, serving as a unique tool to characterize the underlying spacetime structure. $(l,m)=(3,1)$ modes exhibit sharper resonance, which arises within a sequence of successive resonances [10].
@@ II.3 Ambiguities in the definition of excitation factors
The definition of excitation factors is not unique and involves several conventional choices. Throughout this paper, as a reference for the QNM frequencies and excitation factors, we adopt a publicly available dataset [10, 11], hereafter referred to as the Reference Dataset (RD). The RD provides, for $0\leq a/M\leq 1-10^{-6}$, QNM frequencies, angular separation constants, and renormalized angular momenta in the Mano-Suzuki-Takasugi (MST) formalism [46, 47, 32] to 40-digit precision, as well as excitation factors to 16-digit precision. All these quantities are available for gravitational, electromagnetic, and scalar perturbations, while in this paper we focus on gravitational waves. All figures presented in this paper that involve QNM frequencies or excitation factors are based on the RD, either directly or for comparison.
@@
In particular, the integration constant of the tortoise coordinate introduces an inherent ambiguity in the excitation factors. This makes the choice a subtle but important one, and it can sometimes be a source of confusion in practice. If we shift the tortoise coordinate by a constant $C$ as $r_{*}\to r_{*}+C$, the asymptotic amplitudes in Eq. (5) are modified as
@@
Oshita [5] - $h$ $2M=1$ $\frac{1}{2}\ln(1-a^{2}/M^{2})$ HeunG $\Lambda\to 0$ extrapolation
@@
This work - $\Psi_{4}$ $2M=1$ 0 HeunC
@@
Related to these ambiguities, earlier works on computing excitation factors reported different values [55, 9, 5]. Table 1 summarizes this comparison, where we do not include the excitation factors provided in Ref. [55] since they differ from the RD in a way that could not be traced back to a specific convention. As pointed out in Ref. [10], the excitation factors obtained in Ref. [9] included an erroneous factor $e^{i\omega(1-\sqrt{1-a^{2}/M^{2}})}$ inherited from the earlier version of Ref. [32]; correcting this yields consistency with the RD. On the other hand, the excitation factors in Ref. [5] differ by a factor of $(1-a^{2}/M^{2})^{i\omega}$. This likely indicates that the tortoise coordinate employed there involved a corresponding shift of the integration constant, $\tfrac{1}{2}\ln(1-a^{2}/M^{2})$, relative to the conventional one (7)—a shift that diverges in the high-spin limit. That shift coincides with the one later introduced in Ref. [30] to reconcile Eqs. (13) and (15) for the localized source [or more explicitly, Eqs. (24) and (25) below], interpreted as a “start time” of ringdown (see also Sec. 4.2.3 in Ref. [3]).

### v2

@@ Abstract
We investigate how resonant excitation near exceptional points manifests in Kerr black hole ringdown waveforms and examine its extraction. Using waveforms generated by localized initial data, where quasinormal mode amplitudes are given solely by excitation factors, we establish a controlled benchmark for overtone extraction. Applying an iterative fitting method with mirror modes, we analyze a mild resonance in the $(l,m)=(2,2)$ multipole and a sharp resonance in the $(3,1)$ multipole occurring as part of a sequence of successive resonances. For $(2,2)$, we extract the fundamental mode, the first three overtones, and the fundamental mirror mode with relative errors below $10\%$, and show that residual waveforms exhibit the expected damped sinusoids together with distinctive resonance signatures. For $(3,1)$, we demonstrate that resonances can not only amplify but also reduce quasinormal mode excitations, reshaping the overtone hierarchy and rendering the sharp resonance more pronounced in ringdown. Our results clarify the imprint of resonance in Kerr ringdown and highlight both the robustness and limitations of current extraction techniques, providing a foundation for more reliable extraction of higher overtones and for applications to observational data analysis.
@@ I Introduction
The ringdown phase of gravitational waves emitted by perturbed black holes encodes fundamental information about the spacetime geometry in the strong-field regime [1, 2, 3] (see reviews [4, 5, 6]). In this phase, the signal is well described by a superposition of quasinormal modes (QNMs), whose complex frequencies are determined solely by the background spacetime. In addition to the QNM frequencies, their complex amplitudes also carry key information about the strong-field physics, such as the excitation process and the spacetime structure, because the amplitudes can be factorized into a term dependent on the initial data and an excitation factor, the latter of which depends only on the background spacetime.
@@
A particularly intriguing feature arises when QNM frequencies approach each other in the complex plane. This leads to an avoided crossing in which the QNM frequencies trace a hyperbolic trajectory, together with resonant amplification of the corresponding excitation factors in a lemniscate pattern [13, 14], which has been proposed recently as a resolution of the longstanding mystery of the “dissonance” anomaly in Kerr QNM spectrum [15], as well as a novel diagnostic of black hole structure and a potential probe of new physics. Furthermore, it has been derived in Ref. [13] that this phenomenon universally occurs near exceptional points [16], a characteristic feature of non-Hermitian systems that were identified in gravity for the first time. Similar behavior has also been reported in Einstein-Maxwell-axion system [17], reinforcing the potential of the resonance to probe new physics.
@@
The aim of the present work is to identify and characterize the features of resonant excitation in the Kerr ringdown waveform. Our analysis combines three key ingredients: recently computed excitation factors [13, 14], the iterative extraction method [29], and a Kerr ringdown waveform generated by localized initial data [33, 7]. In this waveform, modeled as a delta-function source term, the QNM amplitudes are given solely by the excitation factors. This setup provides a controlled and idealized environment for exploring resonance features and testing their extraction from a waveform composed of an infinite QNM sum and power-law tails. While the response of a localized source has been studied in previous work [33, 7, 34, 18], to our knowledge this is the first time that the fitting of the resulting waveform is explicitly performed. Since QNM amplitudes in more realistic waveforms are also largely governed by excitation factors, the insights gained from this analysis can be directly applied to the fitting of observational signals.
@@ II.1 Teukolsky formalism
Gravitational waves emitted in the ringdown phase of binary black hole mergers are described by linear perturbations on a background black hole spacetime (see [35, 36] for reviews). The perturbations on the Kerr spacetime are governed by a master equation called the Teukolsky equation [37, 38, 39] derived within the Newman-Penrose formalism [40]. The master variable for the Teukolsky equation is $\rho^{-4}\Psi_{4}$, where $\rho=(r-ia\cos\theta)^{-1}$ and $\Psi_{4}$ is a component of the Weyl tensor projected onto null tetrads in the Newman-Penrose formalism. In the asymptotic region far from the perturbed Kerr black hole, $\Psi_{4}$ is related to the gravitational wave strain by $\Psi_{4}=\frac{1}{2}\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}(h_{+}-ih_{\times})$ at infinity [35]. Decomposing the master variable by the Laplace transform
@@
$\int_{0}^{\pi}S^{*}_{lm}(\theta;c)S_{lm}(\theta;c)\sin\theta\mathrm{d}\theta=1,$ (4)
with $c=a\omega$.
@@
and the “up” solution $R^{\text{up}}_{lm}$ satisfying
@@
where $B^{\text{inc}}_{lm}$, $B^{\text{ref}}_{lm}$, $C^{\text{inc}}_{lm}$, and $C^{\text{ref}}_{lm}$ are complex asymptotic amplitudes and $k=\omega-am/(2Mr_{+})$. $r_{*}$ is the tortoise coordinate which satisfies $\mathrm{d}r_{*}/\mathrm{d}r=(r^{2}+a^{2})/\Delta$. We fix the integration constant and employ the conventional tortoise coordinate,
@@
$\Psi_{4}(t,r\to\infty,\theta,\phi)=-\frac{1}{2\pi}\frac{1}{r}\sum_{l=2}^{\infty}\sum_{m=-l}^{l}\int^{\infty}_{-\infty}\mathrm{d}\omega\frac{B^{\text{ref}}_{lm}}{2i\omega B^{\text{inc}}_{lm}}I_{lm}S_{lm}(\theta;a\omega)\frac{e^{im\varphi}}{\sqrt{2\pi}}e^{-i\omega(t-r_{*})}$ (13)
@@ II.2 QNM decomposition
While Eq. (13) involves an integration along the real frequency axis, by enclosing the integration contour on the complex frequency plane, we can rewrite it as a superposition of QNMs together with a power-law tail, provided that no pole arises in the source-dependent term (see Fig. 1). By evaluating the complex frequency integral using the residue theorem, we obtain
@@
where $B_{lmn}$ is the excitation factor [42, 43, 7] defined by
@@
Figure 1: Integration contour for Eq. (13). The cross markers and the jagged line denote the QNM poles and the branch cut, respectively. The integral along the real frequency axis in Eq. (13) is equivalent to closing the contour in the lower half-plane, provided that the contributions from the branch cut and the semicircular arc at infinity are properly taken into account. By the residue theorem, the integral over the closed contour yields the sum over QNMs, while the integral along the branch cut gives rise to the power-law tail. The contribution from the semicircular arc vanishes by Jordan’s lemma.
@@
(a) QNM frequencies for $(l,m)=(2,2)$
(b) Excitation factor for $(l,m)=(2,2)$
(c) QNM frequencies for $(l,m)=(3,1)$
(d) Excitation factor for $(l,m)=(3,1)$
Figure 2: QNM frequencies (left column) and excitation factors (right column) for the $(l,m)=(2,2)$ mode (top row) and $(l,m)=(3,1)$ mode (bottom row) with spin parameter $0\leq a/M\leq 1-10^{-6}$. The cross marker in the left column panels denotes the Schwarzschild value. For $(l,m)=(2,2)$, the range $0.85\leq a/M\leq 0.95$, which corresponds to the repulsion between $n=+5$ and +6+6, is highlighted with thick curves, and the arrow in the right panel shows the direction of increasing spin. For $(l,m)=(3,1)$, the ranges $0.952\leq a/M\leq 0.96$ and $0.972\leq a/M\leq 0.99725$, which correspond to the repulsion between $n=+4$ and +5+5 and between $n=+5$ and +6+6 respectively, are highlighted. All figures involving QNM frequencies or excitation factors use the reference dataset (RD) [13, 14]. Animations of these figures are available in Ref. [51].
Throughout this paper, we adopt a convention where positive overtone indices $n=+0,+1,+2,\cdots$ correspond to ordinary modes, and negative indices $n=-0,-1,-2,\cdots$ to mirror modes. Note that we explicitly include $\pm 0$ to distinguish the original fundamental mode from its mirror counterpart.
@@
There exists a useful symmetry relation between the QNM frequencies and excitation factors of the original and mirror modes. The homogeneous radial Teukolsky equation is invariant under the combined operation of complex conjugation and the replacements $m\to-m$ and $\omega\to-\omega^{*}$. This symmetry implies that the mirror mode frequency is related to the ordinary mode frequency via
@@
Since the same symmetry applies to the radial Teukolsky solutions, the “in” solution satisfies $R^{\text{in*}}_{lm}(\omega)=R^{\text{in}}_{l-m}(-\omega^{*})$. Consequently, the asymptotic amplitudes also satisfy a similar relation, $B^{\text{ref/inc*}}_{lm}(\omega)=B^{\text{ref/inc}}_{l-m}(-\omega^{*})$. Using this relation with Eqs. (16) and (18), the excitation factor of the mirror mode is related to that of the ordinary mode via  [^2] Equation (19) exactly matches Eq. (39) of Ref. [52] aside from notation, while it is inconsistent with Eq. (7) of Ref. [18], which can only be obtained by assuming the opposite sign on the right-hand side of Eq. (19).
@@
Figure 2 shows the QNM frequencies and excitation factors for $(l,m)=(2,2)$ and $(l,m)=(3,1)$ up to the seventh overtone. Throughout this paper, as a reference for the QNM frequencies and excitation factors, we adopt a publicly available dataset [13, 14], hereafter referred to as the reference dataset (RD). For $(l,m)=(2,2)$, we highlight the region with the spin parameter $0.85\leq a/M\leq 0.95$ using thick curves where the avoided crossing and resonant excitation occur. As seen in Fig. 2, the excitation factors of the fifth and sixth overtones exhibit a significant enhancement around $a/M=0.9$, where the corresponding QNM frequencies approach each other. This resonant excitation phenomenon universally occurs in the QNMs in the vicinity of exceptional points [13]. Near this spin value, the excitation factors exhibit a distinctive knot-shaped loop, serving as a unique tool to characterize the underlying spacetime structure.
The $(l,m)=(3,1)$ modes show a complex structure characterized by multiple avoided crossings and resonances [13]. Within this sequence of successive resonances, the resonance between $n=+5$ and +6+6 in the spin range $0.972\leq a/M\leq 0.99725$ stands out as a sharp resonance. In Fig. 2, in addition to this sharp resonance, we also highlight the range $0.952\leq a/M\leq 0.96$ for the mild resonance between $n=+4$ and +5+5 as one of the successive resonances. We shall discuss the role of the mild resonance in the extraction procedure in Sec. IV.4.
Finally, we highlight an important property of the excitation factors in the near-extremal regime. In the extremal limit $a/M\to 1$, QNMs whose frequencies approach the accumulation point ($2M\omega=m$) have excitation factors that tend to zero [53, 7], implying that modes with zero damping rates are not excited. The RD provides high-precision numerical evidence of this behavior at an unprecedented proximity to extremality, up to $a/M=1-10^{-6}$. For instance, the $(l,m,n)=(2,2,+6)$ mode, whose frequency approaches the accumulation point, already shows a nearly vanishing excitation factor at this spin value, whereas the $(2,2,+5)$ mode, which does not approach the accumulation point, retains a finite excitation factor. This evidence strongly suggests that undamped QNMs are not emitted, offering insights into the open problem of the stability of extremal Kerr black holes [54]. It also suggests that in the extremal limit, only a restricted subset of modes—such as the $(2,2,+5)$ mode—retain finite excitation, while others are not excited. This selective excitation raises the possibility that such modes play a special role in the ringdown of extremal Kerr black holes, an intriguing direction for future work on their stability and dynamics.
@@ II.3 Ambiguities in the definition of excitation factors
The definition of excitation factors is not unique and involves several conventional choices. The RD provides, for $0\leq a/M\leq 1-10^{-6}$, QNM frequencies, angular separation constants, and renormalized angular momenta in the Mano-Suzuki-Takasugi (MST) formalism [55, 56, 36] to 40-digit precision, as well as excitation factors to 16-digit precision. All these quantities are available for gravitational, electromagnetic, and scalar perturbations, while in this paper we focus on gravitational waves. All figures presented in this paper that involve QNM frequencies or excitation factors are based on the RD, either directly or for comparison.
@@
In particular, the integration constant of the tortoise coordinate introduces an inherent ambiguity in the excitation factors [34, 52]. This makes the choice a subtle but important one, and it can sometimes be a source of confusion in practice. If we shift the tortoise coordinate by a constant $C$ as $r_{*}\to r_{*}+C$, the asymptotic amplitudes in Eq. (5) are modified as
@@
Related to these ambiguities, earlier works on computing excitation factors reported different values [60, 12, 8]. Table 1 summarizes this comparison, where we do not include the excitation factors provided in Ref. [60] since they differ from the RD in a way that could not be traced back to a specific convention. As pointed out in Ref. [13], the excitation factors obtained in Ref. [12] included an erroneous factor $e^{i\omega(1-\sqrt{1-a^{2}/M^{2}})}$ inherited from the earlier version of Ref. [36]; correcting this yields consistency with the RD. On the other hand, the excitation factors in Ref. [8] differ by a factor of $(1-a^{2}/M^{2})^{i\omega}$. This likely indicates that the tortoise coordinate employed there involved a corresponding shift of the integration constant, $\tfrac{1}{2}\ln(1-a^{2}/M^{2})$, relative to the conventional one (7)—a shift that diverges in the high-spin limit, which limits the reliability of analyses based on the excitation factors of Ref. [8], particularly in the high-spin regime. That shift coincides with the one later introduced in Ref. [34] to reconcile Eqs. (13) and (15) for the localized source [or more explicitly, Eqs. (24) and (25) below], interpreted as a “start time” of ringdown (see also Sec. 4.2.3 in Ref. [6]).
@@
Oshita [8] $h$ $2M=1$ $\frac{1}{2}\ln(1-a^{2}/M^{2})$ HeunG $\Lambda\to 0$ extrapolation
@@
This work $\Psi_{4}$ $2M=1$ 0 HeunC